USE THE FOLLOWING INFORMATION TO ANSWER QUESTIONS 19 -22 The Devlin Corporation, a diversified distribution company, purchases cartons of canned tennis balls from the Questor Company and markets the balls under the Devlin brand name. Devlin started operations on January 1, 2016. In the table below, Quarter 1 represents the time period from January 1, 2016 to March 31, 2016 and Quarter 2 represents the time period from April 1, 2016 to June 30, 2016. Devlin began operations on January 1, 2016 with no inventory Purchases and sales data for Devlins first two quarters of business are shown below: Quarter 1 Sales Purchases 26,000 cartons at $37 per carton January February March 5,000 cartons at $1:2 10,000 cartons at $16 15,000 cartons at $18 Quarter 2 Sales Purchases 63,000 cartons at $45 per cartorn April May June 16,000 cartons at $20 18,000 cartons at $24 26,000 cartons at $30 19. Calculate the Quarter 2 ending inventories (in dollars) using the LIFO inventory valuation method. 20. Calculate Quarter 2 GROSS PROFIT using the LIFO inventory valuation method. 21. Calculate the Quarter 2 ending inventories (in dollars) using the FIFO inventory valuation method. 22. Calculate Quarter 2 GROSS PROFIT using the FIFO inventory valuation method.

PLEASE EXPLAIN ALL STEPS ON HOW YOU GOT TO EACH ANSWER, VERY APPRECIATED!
Answer #1

19: Opening stock in Q2 = Purchases in Q1- Sales in Q1

= (5000+10000+15000)-26000

=4000 units

Total purchases in units= 16000+18000+ 26000 = 60000

Total sales in units= 63000

Closing stock = (Opening stock+ Purchases- Sales) = (4000+60000-63000) = 1000 units

Value based on LIFO = 1000* Jan price

= 1000*12

=$12000

20: Value of opening stock = 4000*12 = 48000

Purchases in Q2 = 16000*20 + 18000*24+ 26000*30 = 1532000

Sales in Q2 = 63000*45=2835000

Closing stock =12000 (As per 19)

Gross Profit = Sales - (Opening stock+ Purchases-Closing stock)

= 2835000- (48000+1532000-12000)

=1267000

21: Opening stock in Q2 = Purchases in Q1- Sales in Q1

= (5000+10000+15000)-26000

=4000 units

Total purchases in units= 16000+18000+ 26000 = 60000

Total sales in units= 63000

Closing stock = (Opening stock+ Purchases- Sales) = (4000+60000-63000) = 1000 units

Value based on FIFO = 1000* June price

= 1000*30

=$30000

22: Value of opening stock = 4000*18 = 72000

Purchases in Q2 = 16000*20 + 18000*24+ 26000*30 = 1532000

Sales in Q2 = 63000*45=2835000

Closing stock =30000 (As per Q21)

Gross Profit = Sales - (Opening stock+ Purchases-Closing stock)

= 2835000- (72000+1532000-30000)

=1261000